# HaluMem: Evaluating Hallucinations in Memory Systems of Agents

### v1

@@ 1 Introduction
A memory system serves as the fundamental infrastructure for organizing and managing information based on the history of human–AI conversations. It extracts, structures, and continuously updates key information generated across multi-turn interactions between users and AI systems, retrieving and injecting this information into the model as needed to support personalization and long-term consistency [li2025memos_long, memobase, kang2025memory, rasmussen2025zep, supermemory]. Specifically, a memory system identifies stable user profiles, narratives, and events from dialogues and stores them as plaintext entries enriched with metadata. When new queries or tasks arise, the system retrieves and selectively integrates relevant memories based on the current intent and context, enabling the AI system to “remember and correctly utilize” user information, thereby preserving semantic coherence, behavioral consistency, and preference alignment. Representative systems such as MemOS [li2025memos_long], Mem0 [chhikara2025mem0], and Memobase [memobase] continuously record user profiles, events, and evolving preferences, supporting the creation, revision, and tracking of memories to construct a system-level memory layer with structured management capabilities.
@@ 2.1 Memory System
RAG [lewis2020rag] Plain Text CUD Good No
GraphRAG [edge2025graphrag] Plain Text CUD fair Yes
Supermemory [supermemory] Plain Text CUDE fair–Excellent Yes
@@
Table 1: Memory tools, ranging from early memory plug-ins to advanced memory systems.
@@
As shown in Table 1, various forms of external memory modules have been proposed to address the limitations of parameterized memory. Early external memory mechanisms were primarily represented by Retrieval-Augmented Generation (RAG). In particular, RAG [lewis2020rag] introduces an external plaintext knowledge retrieval mechanism. Before generation, relevant documents are retrieved from a vector database and incorporated into the model’s input, enabling controllable and updatable external memory. This approach offers high manageability, and because the external memory is transparent and editable, it exhibits a relatively low degree of memory hallucination. However, traditional RAG systems primarily rely on text-based memory structures and lack explicit modeling of inter-entity relationships, meaning they do not support graph structures. Consequently, they remain limited in handling complex knowledge reasoning and maintaining long-term consistency. Building upon this, GraphRAG [edge2025graphrag] further integrates a knowledge graph structure, organizing and retrieving knowledge in the form of entity–relation pairs. By leveraging graph indexing and multi-hop path retrieval, GraphRAG significantly enhances the representational capacity and retrieval accuracy of structured knowledge, leading to improved performance in relational reasoning. Nevertheless, the construction and maintenance of graph structures entail high costs, and synchronization during updates introduces additional complexity. As a result, GraphRAG demonstrates moderate manageability, inferior to that of RAG, and may introduce additional memory hallucinations due to inconsistent updates among graph nodes or edges.
@@
With the growing demand for AI systems capable of personalized interaction and long-term learning, researchers have begun exploring memory systems that possess genuine long-term maintainability and operational controllability. Memobase [memobase] primarily targets user-level long-term memory scenarios. This framework employs a plaintext structure with metadata (such as timestamps and data sources) to record user preferences and interaction histories. During user interactions, the system can access current user profiles and recent events via an API to generate context snippets, which are injected into the model’s prompt, enabling the model to “remember” the user. This design performs well in managing user profiles and maintaining event timelines, although certain risks of hallucination may arise during the memory extraction process. Mem0 [chhikara2025mem0] adopts a metadata-enriched plaintext storage format with fields such as id, timestamp, source, ttl (time to live), and version, supporting a wider range of memory operations—Create/Extract, Update, Delete, and Expand/Enrich (CUDE). For example, it implements conflict detection and memory merging mechanisms to ensure consistency and traceability. MemOS attempts to abstract memory as a system-level resource by unifying the management of three types of memory: parametric memory, activation memory, and explicit (plaintext) memory. Through lifecycle control, version management, and migration mechanisms, MemOS [li2025memos_long] enables cross-model and cross-session memory sharing and integration. However, although the introduction of graph structures enhances the representational expressiveness among memories, it simultaneously increases management complexity, such as maintaining update consistency. Moreover, compared with unstructured memory, graph-based memory is more susceptible to hallucination during both generation and management.
@@ 2.2 Evaluation Hallucinations in Memory Systems
Generation hallucinations are the most extensively studied type of hallucination in current research and are typically divided into two categories: factual hallucinations and faithfulness hallucinations [huang2025survey]. Factual hallucinations assess whether the model output aligns with objective facts, whereas faithfulness hallucinations evaluate whether the output remains faithful to the given context or source information. Around these two categories, researchers have proposed a variety of mature evaluation methods and metric systems, including: external-retrieval-based factual verification, which compares generated content with external knowledge bases; model-internal-state-based reliability assessment, which analyzes attention distributions or activation patterns to estimate hallucination risk; behavior-based evaluation of output consistency and verifiability; and uncertainty- or LLM-discriminator-based automated hallucination detection. These approaches have substantially improved the interpretability and quantifiability of hallucinations at the generation level, leading to relatively mature and systematic detection frameworks for generation hallucinations.
In contrast, research on memory hallucinations remains at an early stage. Representative studies include LoCoMo, LongMemEval, PrefEval, and PersonaMem. LoCoMo [maharana2024evaluating] is one of the earliest benchmarks focusing on long-context memory capabilities. It primarily evaluates the model’s accuracy in retaining and recalling information within ultra-long texts or multi-turn dialogues. Its task design includes factual recall, entity tracking, and event sequence reasoning to measure the degree of contextual forgetting and confusion. LongMemEval [wulongmemeval] emphasizes the accuracy of information retention and retrieval in multi-turn interactions. This benchmark introduces two core metrics—Information Retention Rate and Memory Recall Accuracy—to evaluate the model’s consistency of knowledge across different temporal spans. PrefEval [zhao2025do] focuses on assessing a model’s ability to adhere to user preferences throughout long-term dialogue. It constructs preference–query pairs and inserts numerous irrelevant dialogues as distractors to test the model’s preference inference, long-context retrieval, and personalization proactiveness. PersonaMem [jiang2025know] evaluates identity consistency and memory tracking ability by simulating user personas and event histories. It introduces several evaluation dimensions, including Persona Consistency, Traceability, and Update Accuracy, targeting the personalized aspects of long-term memory evaluation.
@@ 4 Methodology for Constructing HaluMem
Stage 1: Persona Construction. This stage initiates the HaluMem dataset construction by creating virtual users with complete persona profiles to simulate real participants in later human–AI dialogues. Each persona includes three parts: Core Profile Information, Dynamic State Information, and Preference Information. The core profile captures stable background traits; the dynamic state reflects current circumstances such as career, health, and relationships; and the preferences define personal tastes. Each user receives six to eight randomly assigned preferences across areas like food, music, and film. The core profile provides a static foundation, while the dynamic and preference elements, which evolve randomly, add realism, diversity, and rich material for memory extraction. An initial timestamp ensures that all personas reflect a consistent starting point in time. To enhance the authenticity of virtual users, user seeds are randomly sampled from the Persona Hub [^3] A collection of one billion diverse personas automatically curated from web data [tao2024personahub], and rule‑based procedures are applied to generate structured persona drafts. GPT‑4o then verifies and refines them. See Appendix D.1 for an example.
@@
Within this framework, career events serve as the narrative backbone, while init and daily events provide necessary background and contextual details. Through the integration and chronological alignment of the three event types, this stage produces a coherent and complete event sequence that functions as the user’s memory transaction log. See Appendix D.2 for event examples.
@@
Stage 4: Session Summaries and Memory Points. This stage transforms the structured event flow from Stage 3 into realistic session summaries and detailed memory points. For each event, we create a human–AI dialogue scenario shaped by the user’s motivation. The system has access to the current persona profile, along with all prior events and memory points, ensuring logical, causal, and consistent generation. As events unfold, the persona profile is dynamically updated to reflect the user’s evolving state. Each memory point includes its content, type(persona, event, or relationship), and importance, with updated entries preserving replaced information for traceability. More details provided in Appendices A.1 and D.3.
@@
Stage 5: Session Generation. This stage converts the structured event flow and memory points from previous stages into complete, multi-turn dialogues that are context-rich, goal-driven, and adversarially challenging. The process has three steps: adversarial content injection, multi-turn dialogue generation, and memory self-verification. Adversarial content injection adds distractor memories [^4] False but similar memories that the AI naturally uses while the user stays silent, mimicking realistic information contamination. Memory self-verification checks and refines each memory point for consistency with the generated dialogues. Overall, this stage simulates how memory is formed, maintained, and challenged in realistic conversations, producing data that test long-term memory performance and hallucination resistance. Examples appear in Appendix D.3.
@@
Stage 6: Question Generation. The final stage constructs a set of memory-related question–answer pairs based on the sessions and memory points generated previously. Six categories of memory evaluation questions are predefined, and the number and types of questions are programmatically allocated according to event type and complexity to ensure balanced coverage. For each career event, all its sub-stages are integrated into a single unit to increase reasoning depth and complexity. Each question–answer pair is annotated with a difficulty level and accompanied by traceable evidence, explicitly linking the answer to the supporting memory points. See Appendices A.2 and D.3 for details.
@@
Human Annotation. To verify the quality of HaluMem, we conducted human annotation on part of the sessions in HaluMem-Medium, covering both memory points and question–answer pairs. For each user, 35 sessions were randomly selected, totaling 700 sessions (over 50% of the dataset). Eight annotators with at least a bachelor’s degree rated each session on Correctness, Relevance, and Consistency. After 10 days of annotation, the results showed a correctness rate of 95.70%, an average relevance score of 9.58, and an average consistency score of 9.45. These results demonstrate the high quality and reliability of the HaluMem benchmark. More details are provided in Appendix B.
@@ 5.1 Memory Extraction
Memory Completeness (Anti-Amnesia)
@@ 6.1 Experimental Setup
We conducted a comprehensive evaluation of several state-of-the-art memory systems on HaluMem, including Mem0 (both standard and graph versions) [chhikara2025mem0], Supermemory [supermemory], and Memobase [memobase]. [^7] Due to the large scale of HaluMem and the significantly slower evaluation speed of Zep [rasmussen2025zep], we plan to include its evaluation results in future work. Each memory system was independently evaluated in two subsets, HaluMem-Medium and HaluMem-Long, with efforts made to ensure consistent parameter configurations across evaluations.
To automate the evaluation of three core tasks, memory extraction, memory updating, and memory question answering, we use GPT-4o for consistency determination and scoring. We designed various prompt templates to guide the automated evaluation by GPT-4o (See Appendix C.2 for details.). In the memory updating task, we retrieved the 10 most relevant memories from the memory system for each memory labeled as "update type" for verification. In the memory question answering task, we retrieved 20 most relevant memories for each question to assist in generating answers, using GPT-4o uniformly as the answer generation model. The prompt templates used for answer generation across different memory systems are provided in Appendix C.1.
@@ 6.2.1 Overall Evaluation on HaluMem
Overall, almost all memory systems perform worse on HaluMem-Long than on HaluMem-Medium, with Mem0, Mem0-Graph, and Memobase showing particularly notable declines. Notably, except for Supermemory, all other systems extract significantly fewer memories on HaluMem-Long than on HaluMem-Medium. This indicates that future memory systems need to improve their ability to process irrelevant information and distinguish between high- and low-value memories.
@@
In the memory updating task, all systems achieve correct update rates below 26%, and except for Supermemory, their performance drops considerably on HaluMem-Long. Systems showing better performance in memory integrity also tend to exhibit higher update accuracy, but all suffer omission rates above 50%. This issue primarily stems from insufficient coverage in memory extraction: when the pre-update memories are not extracted, related updates cannot be properly processed. Moreover, the fact that all systems exhibit hallucination rates below 1% does not necessarily imply strong hallucination suppression, since very few samples actually enter the update stage. Overall, current systems face a clear bottleneck in memory updating: the extraction and updating stages lack stable linkage, resulting in low accuracy and high omission rates.
In the memory question-answering task, the best-performing systems are also those achieve the highest memory integrity, further highlighting the crucial role of memory extraction. For example, Mem0 and Mem0-Graph show clear performance declines on HaluMem-Long compared to HaluMem-Medium, which strongly correlates with their substantial reduction in extracted memory points. However, all systems achieve answer accuracies below 55%, with both hallucination rate and omission rate remaining high, and their overall performance further decreases on HaluMem-Long. This demonstrates that current memory systems’ QA performance depends heavily on the sufficiency and accuracy of upstream memory extraction, and remains prone to factual deviation and memory confusion under interference or extended context conditions.

### v2

@@ 1 Introduction
A memory system serves as the fundamental infrastructure for organizing and managing information based on the history of human–AI conversations. It extracts, structures, and continuously updates key information generated across multi-turn interactions between users and AI systems, retrieving and injecting this information into the model as needed to support personalization and long-term consistency [li2025memos_long, memobase, kang2025memory, rasmussen2025zep, supermemory]. Specifically, a memory system identifies stable user profiles, narratives, and events from dialogues and stores them as plaintext entries enriched with metadata. When new queries or tasks arise, the system retrieves and selectively integrates relevant memories based on the current intent and context, enabling the AI system to “remember and correctly utilize” user information, thereby preserving semantic coherence, behavioral consistency, and preference alignment. Representative systems such as MemOS [li2025memos_long], Mem0 [chhikara2025mem0], Zep [rasmussen2025zep], Supermemory [supermemory], and Memobase [memobase] continuously record user profiles, events, and evolving preferences, supporting the creation, revision, and tracking of memories to construct a system-level memory layer with structured management capabilities.
@@ 2.1 Memory System
Supermemory [supermemory] Plain Text (with Metadata) CUDE fair–Excellent Yes
@@
Various forms of external memory modules have been proposed to address the limitations of parameterized memory. Early external memory mechanisms were primarily represented by Retrieval-Augmented Generation (RAG). In particular, RAG [lewis2020rag] introduces an external plaintext knowledge retrieval mechanism. Before generation, relevant documents are retrieved from a vector database and incorporated into the model’s input, enabling controllable and updatable external memory. This approach offers high manageability, and because the external memory is transparent and editable, it exhibits a relatively low degree of memory hallucination. However, traditional RAG systems primarily rely on text-based memory structures and lack explicit modeling of inter-entity relationships, meaning they do not support graph structures. Consequently, they remain limited in handling complex knowledge reasoning and maintaining long-term consistency. Building upon this, GraphRAG [edge2025graphrag] further integrates a knowledge graph structure, organizing and retrieving knowledge in the form of entity–relation pairs. By leveraging graph indexing and multi-hop path retrieval, GraphRAG significantly enhances the representational capacity and retrieval accuracy of structured knowledge, leading to improved performance in relational reasoning. Nevertheless, the construction and maintenance of graph structures entail high costs, and synchronization during updates introduces additional complexity. As a result, GraphRAG demonstrates moderate manageability, inferior to that of RAG, and may introduce additional memory hallucinations due to inconsistent updates among graph nodes or edges.
@@
With the growing demand for AI systems capable of personalized interaction and long-term learning, researchers have begun exploring memory systems that possess genuine long-term maintainability and operational controllability, as summarized in Table 1. Supermemory [supermemory] provides long-term memory for language models by combining document retrieval and user-specific memory. It integrates both retrieval-augmented generation and agent memory, using a contextual graph that captures temporal, relational, and personal information, enabling consistent and personalized responses across interactions. Memobase [memobase] focuses on user-level long-term memory by recording preferences and interaction histories in a plaintext structure. During interactions, it dynamically generates context snippets from user profiles and recent events to enable personalized recall, though some risk of hallucination may occur during memory extraction. Zep [rasmussen2025zep] introduces a context engineering framework that integrates agent memory, Graph RAG, and context assembly capabilities, with its core component Graphiti enabling temporally-aware synthesis of conversational and business data for personalized long-term context. Mem0 [chhikara2025mem0] employs a metadata-enriched plaintext storage format that supports comprehensive memory operations (Create/Extract, Update, Delete, and Expand/Enrich), incorporating conflict detection and memory merging to ensure consistency and traceability. MemOS [li2025memos_long] attempts to abstract memory as a system-level resource by unifying the management of three types of memory: parametric memory, activation memory, and explicit (plaintext) memory. Through lifecycle control, version management, and migration mechanisms, MemOS enables cross-model and cross-session memory sharing and integration. However, while graph structures enhance the expressiveness of memory representation, they also increase management complexity and make the system more prone to hallucination.
@@ 2.2 Evaluation Hallucinations in Memory Systems
  • * “Several years” for PersonaMem is inferred from the paper and dataset, not explicitly labeled.
@@
Generation hallucinations are the most extensively studied type of hallucination in current research and are typically divided into two categories: factual hallucinations and faithfulness hallucinations [huang2025survey]. Factual hallucinations assess whether the model output aligns with objective facts, whereas faithfulness hallucinations evaluate whether the output remains faithful to the given context or source information. Around these two categories, researchers have proposed a variety of mature evaluation methods and metric systems, including: external-retrieval-based factual verification [DBLP:journals/corr/abs-2502-13622], which compares generated content with external knowledge bases; model-internal-state-based reliability assessment [su-etal-2024-unsupervised, DBLP:conf/iclr/0026L0GWTFY24], which analyzes attention distributions or activation patterns to estimate hallucination risk; behavior-based evaluation of output consistency and verifiability [liang-etal-2024-learning]; and uncertainty- or LLM-discriminator-based automated hallucination detection [kang2025uncertaintyquantificationhallucinationdetection, DBLP:journals/corr/abs-2505-08200]. These approaches have substantially improved the interpretability and quantifiability of hallucinations at the generation level, leading to relatively mature and systematic detection frameworks for generation hallucinations.
In contrast, research on memory hallucinations remains at an early stage. Representative studies include LoCoMo, LongMemEval, PrefEval, and PersonaMem (Table 2). The early benchmark LoCoMo [maharana2024evaluating] focuses on memory retention under long-context settings. It adopts an end-to-end evaluation paradigm, assessing models through question answering, summarization, and generation tasks to test factual recall and event tracking over ultra-long texts. Although the dataset is relatively large (about 7.5k questions), it only covers a time span of several months and lacks an explicit memory updating mechanism, primarily reflecting the model’s capability for static information retention. Subsequently, LongMemEval [wulongmemeval] extends this framework by introducing metrics such as Information Retention Rate and Memory Recall Accuracy, covering approximately 2.5 years of multi-turn dialogue and incorporating explicit memory updates to quantify knowledge consistency across time. This represents a shift from static evaluation toward dynamic memory modeling. In the direction of personalization, PrefEval [zhao2025do] evaluates a model’s ability to maintain and follow user preferences over long-term interactions, using generation and classification tasks to assess preference consistency, though it remains limited to persona-level memory. PersonaMem [jiang2025know] further constructs simulated user personas and event histories, employing multiple-choice evaluations to assess persona consistency, traceability, and update accuracy. With a longer time span (on the order of years), it provides a more representative benchmark for personalized long-term memory assessment.
@@ 4 Methodology for Constructing HaluMem
Stage 1: Persona Construction. This stage initiates the HaluMem dataset construction by creating virtual users with complete persona profiles to simulate real participants in later human–AI dialogues. Each persona includes three parts: Core Profile Information, Dynamic State Information, and Preference Information. The core profile captures stable background traits; the dynamic state reflects current circumstances such as career, health, and relationships; and the preferences define personal tastes. Each user receives six to eight randomly assigned preferences across areas like food, music, and film. The core profile provides a static foundation, while the dynamic and preference elements, which evolve randomly, add realism, diversity, and rich material for memory extraction. An initial timestamp ensures that all personas reflect a consistent starting point in time. To enhance the authenticity of virtual users, user seeds are randomly sampled from the Persona Hub [^3] A collection of one billion diverse personas automatically curated from web data [tao2024personahub], and rule‑based procedures are applied to generate structured persona drafts. GPT‑4o then verifies and refines them. See Appendix E.1 for an example.
@@
Within this framework, career events serve as the narrative backbone, while init and daily events provide necessary background and contextual details. Through the integration and chronological alignment of the three event types, this stage produces a coherent and complete event sequence that functions as the user’s memory transaction log. See Appendix E.2 for event examples.
@@
Stage 4: Session Summaries and Memory Points. This stage transforms the structured event flow from Stage 3 into realistic session summaries and detailed memory points. For each event, we create a human–AI dialogue scenario shaped by the user’s motivation. The system has access to the current persona profile, along with all prior events and memory points, ensuring logical, causal, and consistent generation. As events unfold, the persona profile is dynamically updated to reflect the user’s evolving state. Each memory point includes its content, type(persona, event, or relationship), and importance, with updated entries preserving replaced information for traceability. More details provided in Appendices A.1 and E.3.
@@
Stage 5: Session Generation. This stage converts the structured event flow and memory points from previous stages into complete, multi-turn dialogues that are context-rich, goal-driven, and adversarially challenging. The process has three steps: adversarial content injection, multi-turn dialogue generation, and memory self-verification. Adversarial content injection adds distractor memories [^4] False but similar memories that the AI naturally uses while the user stays silent, mimicking realistic information contamination. Memory self-verification checks and refines each memory point for consistency with the generated dialogues. Overall, this stage simulates how memory is formed, maintained, and challenged in realistic conversations, producing data that test long-term memory performance and hallucination resistance. Examples appear in Appendix E.3.
@@
Stage 6: Question Generation. The final stage constructs a set of memory-related question–answer pairs based on the sessions and memory points generated previously. Six categories of memory evaluation questions are predefined, and the number and types of questions are programmatically allocated according to event type and complexity to ensure balanced coverage. For each career event, all its sub-stages are integrated into a single unit to increase reasoning depth and complexity. Each question–answer pair is annotated with a difficulty level and accompanied by traceable evidence, explicitly linking the answer to the supporting memory points. See Appendices A.2 and E.3 for details.
@@
Human Annotation. To verify the quality of HaluMem, we conducted human annotation on part of the sessions in HaluMem-Medium, covering both memory points and question–answer pairs. For each user, 35 sessions were randomly selected, totaling 700 sessions (over 50% of the dataset). Eight annotators with at least a bachelor’s degree rated each session on Correctness, Relevance, and Consistency. After 10 days of annotation, the results showed a correctness rate of 95.70%, an average relevance score of 9.58, and an average consistency score of 9.45. These results demonstrate the high quality and reliability of the HaluMem benchmark. More details are provided in Appendix C.
@@ 5.1 Memory Extraction
Memory Integrity (Anti-Amnesia)
@@ 6.1 Experimental Setup
We conducted a comprehensive evaluation of several state-of-the-art memory systems on HaluMem, including Mem0 (both standard and graph versions) [chhikara2025mem0], Memobase [memobase], Supermemory [supermemory], and Zep [rasmussen2025zep]. [^7] MemOS will also be included in the evaluation as soon as possible. Each memory system was independently evaluated in two subsets, HaluMem-Medium and HaluMem-Long, with efforts made to ensure consistent parameter configurations across evaluations.
To automate the evaluation of three core tasks, memory extraction, memory updating, and memory question answering, we use GPT-4o for consistency determination and scoring. We designed various prompt templates to guide the automated evaluation by GPT-4o (See Appendix D.2 for details.). In the memory updating task, we retrieved the 10 most relevant memories from the memory system for each memory labeled as "update type" for verification. In the memory question answering task, we retrieved 20 most relevant memories for each question to assist in generating answers, using GPT-4o uniformly as the answer generation model. The prompt templates used for answer generation across different memory systems are provided in Appendix D.1.
Some memory systems required specific Configurations due to their unique interfaces and constraints; details are provided in Appendix B.
@@ 6.2.1 Overall Evaluation on HaluMem
Zep - - - - - 37.35% 0.48% 62.14% 50.19% 22.51% 27.30%
  • Note: since Zep does not provide a Get Dialogue Memory API, it is not possible to calculate Memory Integrity or Memory Accuracy. For details, see Appendix B.
@@
Overall, almost all memory systems perform worse on HaluMem-Long than on HaluMem-Medium, with Mem0, Mem0-Graph, and Memobase showing particularly notable declines. Notably, except for Supermemory and Zep, all other systems extract significantly fewer memories on HaluMem-Long than on HaluMem-Medium. This indicates that future memory systems need to improve their ability to process irrelevant information and distinguish between high- and low-value memories.
@@
In the memory updating task, all systems achieve correct update rates below 50%, and except for Supermemory, their performance drops considerably on HaluMem-Long. Systems showing better performance in memory integrity also tend to exhibit higher update accuracy, but all suffer omission rates above 50%. This issue primarily stems from insufficient coverage in memory extraction: when the pre-update memories are not extracted, related updates cannot be properly processed. Moreover, the fact that all systems exhibit hallucination rates below 2% does not necessarily imply strong hallucination suppression, since very few samples actually enter the update stage. Overall, current systems face a clear bottleneck in memory updating: the extraction and updating stages lack stable linkage, resulting in low accuracy and high omission rates.
In the memory question-answering task, the best-performing systems are also those that perform well in memory integrity and memory updating, further highlighting the crucial role of memory extraction. For example, Mem0 and Mem0-Graph show clear performance declines on HaluMem-Long compared to HaluMem-Medium, which strongly correlates with their substantial reduction in extracted memory points. However, all systems achieve answer accuracies below 56%, with both hallucination rate and omission rate remaining high, and their overall performance further decreases on HaluMem-Long. This demonstrates that current memory systems’ QA performance depends heavily on the sufficiency and accuracy of upstream memory extraction, and remains prone to factual deviation and memory confusion under interference or extended context conditions.